# TerraCodec: Compressing Earth Observations

### v1

@@ Appendix C Latent Repacking and FlexTEC
Scope and notation.
FlexTEC is the flexible–rate variant of TEC–TT, obtained by integrating Latent Repacking and masked training. It uses the same analysis/synthesis (ELIC) backbone and temporal transformer stack as TEC–TT, with latent channel width $d_{\text{lat}}{=}192$. After tokenizing the current frame into T=16T{=}16 tokens per patch, repacking groups channels into $T$ channel-slice tokens (Sec. 3.2). Consequently, FlexTEC exposes 16 discrete quality levels via the token budget $K\in\{1,\dots,16\}$, applied consistently across all patches in an image and frames in a sequence. Each token carries $k=d_{\text{lat}}/T=12$ channels shared across all spatial positions. For all rate–distortion (RD) visualizations in this paper, we report curves for budgets $K=\{1,2,3,4,5,6,7,8,12,16\}$, spanning from the lowest-rate (K=1K{=}1) to the highest-quality (K=16K{=}16) operating points.
Implementation differences vs. TEC–TT.
FlexTEC is architecturally identical to TEC–TT except for: (i) the permutation that repacks tokens (Sec. 3.2); (ii) token masking with a learned mask token $m\in\mathbb{R}^{d_{\text{lat}}}$ used during training to replace dropped tokens; and (iii) the masked-rate objective with budget sampling. All other layers, dimensions, and hyperparameters remain unchanged.
FlexTEC is trained with the same hyperparameters as TEC–TT, but for 400k steps (vs. 300k for TEC–TT) to account for the task’s added complexity. A single checkpoint trained at $\lambda{=}800$ (slightly higher than the TEC–TT maximum of 700) is used to cover the full bitrate range under masked training. Empirically, the $T/K$ scaling in the rate term shifts the effective operating point toward lower rates for the same $\lambda$, motivating this increase.
Objective and inference.
With mask $M$, the rate is computed on unmasked tokens and, to prevent information collapse into the earliest tokens, upweighted by $T/K$:
@@
Masking for variable-rate robustness.
Repacking latents alone is insufficient: a model trained only with full-token inputs learns to rely on all tokens and collapses when some are dropped. We therefore train with masked budgets: sample $K\in\{1,\dots,T\}$, replace the last TKT{-}K tokens by a learned mask vector $m\in\mathbb{R}^{d_{lat}}$, and compute the rate only on unmasked tokens $R(M)$ (as defined in Eq. 3). For stability we use teacher forcing: masking applies only to the current frame’s tokens for the rate term, while the temporal encoder always consumes the real quantized past latents $(\hat{\mathbf{y}}_{i-2},\hat{\mathbf{y}}_{i-1})$. Inspired by Bachmann et al. (2025), budgets are drawn from a categorical distribution biased toward larger values, $\Pr(K{=}k)\propto k$ (i.e., the multiset $\{1,2,2,\dots,T,\dots,T\}$), which trains the model frequently near high-rate operation while still exposing it to low-budget regimes. Masked tokens are replaced by a learnable per-channel vector $m\in\mathbb{R}^{d_{lat}}$ ($d_{lat}{=}192$), initialized uniformly in [1,1][-1,1], and shared across positions. This provides a stable placeholder during training while allowing the image decoder to learn how to interpret missing content.
@@
Inference filling.
At inference, dropped tokens are not transmitted. By default, we fill them with the transformer’s predicted means, i.e., $\hat{\mathbf{y}}_{i,b,t}\leftarrow\mu_{i,b,t}$ for frame $i$, latent block $b$, and token $t$, which leverages the learned temporal prior to improve reconstruction quality at a given bitrate. As a lighter alternative (reduced compute), we can instead substitute the learned mask vector $m$ for all dropped tokens.
Appendix D Evaluation details and baseline methods
Baseline quality settings.
We evaluate classical codecs across the following quality grids:
@@ Appendix D Evaluation details and baseline methods
Implementations and bit depth.
JPEG and WebP are executed via Pillow, supporting only 8-bit input. JPEG2000 is run per band using Glymur/OpenJPEG, which allows for 8- or 16-bit quantization and lossy compression via target ratios. HEVC encoding is performed with FFmpeg and x265, using raw 12-bit monochrome input (pixel format yuv400p12le), CRF=Q=Q, -preset medium, and -tune psnr. Bitstream sizes are measured directly from encoded outputs to compute rate. All codecs operate per band (grayscale), and we report bits-per-pixel-band-frame (bppbf) by aggregating bitstream sizes across all bands and frames.
Rate metric (bppbf).
We report rate as bits per pixel–band–frame (bppbf):
@@
Appendix E Quantitative results
@@
E.1 PSNR range sensitivity
@@
E.2 Temporal conditioning effects
We evaluate how temporal context influences fixed-rate TEC–TT models at inference. To isolate genuine temporal gains from the overhead of early bootstrap frames, we vary the number of available past frames and additionally analyze P-frame performance.
@@ E.2 Temporal conditioning effects
Figure 12 shows corresponding RD curves evaluated on P-frames only—the last two frames in each four-frame sequence, consistent with TEC–TT’s training context of two past frames. For non-temporal codecs, this selection simply aligns the evaluation set with TEC–TT. The performance gap between TEC–TT and image-only codecs (TEC–FP, TEC–ELIC) widens under this evaluation, reflecting the amortized cost of the initial bootstrap frames. When temporal conditioning is disabled and TEC–TT is run in “image mode,” its performance closely follows TEC–ELIC, consistent with their shared analysis/synthesis backbone.
Table 6: Effect of inference context $c$ on TEC–$T$ (trained with 2-frame context, $\lambda{=}5$). We report bits-per-pixel–band–frame (bppbf, $\downarrow$) and PSNR-per-band (65k, $\uparrow$). Context $c\in\{0,1,1{+}1,2\}$ denotes the number of conditioned frames. The P-frames row evaluates only the last two frames (full context) of each 4-frame sequence. Percent changes are relative to c=0c{=}0.
@@
E.3 Flexible-rate behavior
@@
Fig. 13 shows the effect of masking in flexible‐rate training by comparing FlexTEC (Latent Repacking with masking) to an variant trained without masking. The latter deteriorates sharply when tokens are dropped at test time—its RD curve is unstable and substantially below FlexTEC—whereas FlexTEC degrades smoothly and remains roughly parallel to fixed‐rate baselines. This confirms that masking is essential for stable variable‐rate performance.
Fig. 14 compares FlexTEC on P-frames (last two frames, full context under c=2c{=}2) versus all frames. The gap is notably larger than for fixed-rate TEC–TT, particularly at low rates. We hypothesize two compounding effects: (i) latent repacking with masked training encourages FlexTEC to concentrate scene-wide, high-utility content into the earliest tokens of bootstrap frames, lowering their cost relative to fixed-rate models; and (ii) for fully conditioned P-frames, the same curriculum distributes information more evenly across tokens, so truncation retains most of the essentials. Together, these effects yield stronger amortization when excluding bootstrap frames, with the benefit most pronounced in the high-compression regime.
@@
Appendix F Qualitative examples
This section complements the main paper with additional visual examples. We first compare reconstructions across all codecs on representative SSL4EO-S12 samples (Sec. F.1). We then examine model beliefs and forecasting behavior (Sec. F.2) for TEC-TT and FlexTEC. We assesss how dropping tokens impacts TEC-TT and the importance of token masking for Latent Repacking. We also show how TEC–TT forecasts the next frame from past context, contrasting mean predictions with stochastic samples.
@@ Appendix F Qualitative examples
F.1 General reconstructions
We compare TerraCodec-TT (TEC-TT), TerraCodec-FP (TEC-FP), and classical codecs (JPEG2000, WebP) on SSL4EO-S12 v1.1 validation samples at matched rate $\approx 0.20$ bppbf. Each row in Fig. F.1) shows the original image and reconstructions from each codec, annotated with the average bppbf and PSNR-per-band (65k clipped) across the sequence.
F.2 Model beliefs and forecasting
@@
F.2.1 Token budget comparison
Figure 16: Effect of filling dropped tokens (K=5K{=}5) with model prior predictions at inference for different TEC–$T$ variants. Vanilla TEC–$T$ exhibits spatial holes and banding when tokens are removed. Adding Latent Repacking without masking improves quality but leaves uneven detail, while FlexTEC preserves global layout and reduces artifacts.
@@ F.2.1 Token budget comparison
F.2.2 Forecasts from past context
@@
Figure 17: TEC–TT forecasts using only past context ($K=0$). Model beliefs are visualized via the mean prediction ($\mu$) and full prior sampling from $\mathcal{N}(\mu,\sigma^{2})$. The $\mu$-forecast reconstructs coherent large-scale structure, while sampling reveals plausible variations (e.g., clouds, surface texture) that reflect the model’s uncertainty.
@@
We probe TEC–TT’s model beliefs by predicting the current frame from past context only, using either the predicted mean or samples from the distribution (Fig. 17). The $\mu$-forecast reliably captures large-scale structure, while sampling from the full prior (mean and variance) expresses context-aware uncertainty, primarily reflecting cloud variability. This illustrates the learned distribution rather than a single point estimate. The conservative single-point forecast ($\mu$-forecast) produces a clear-sky prediction for the next frame, enabling TEC–TT to perform cloud removal as evaluated in AllClear.
Appendix G Additional details on cloud inpainting
@@
We provide extended results for the AllClear cloud inpainting benchmark (Zhou et al., 2024), complementing Sec. 4.3 and Table 1. TEC–TT is applied without task-specific fine-tuning, leveraging only its latent temporal predictions.
Experimental setup.
@@
We follow the AllClear evaluation protocol, reporting metrics across all spectral bands (not per-band). We compare against heuristic baselines (LeastCloudy, Mosaicing) and zero-shot neural methods (CTGAN, DiffCR, PMAA, U-TILISE, UnCRtainTS). LeastCloudy selects the input image with the lowest cloud+shadow coverage, while Mosaicing fills each pixel by copying a single clear value, averaging if multiple are clear, or using 0.5 if none are clear. On AllClear, these heuristics rank among the top three zero-shot methods, outperforming most neural baselines without fine-tuning.
@@
Besides reporting metrics on the full test set, we also evaluate subsets stratified by cloudiness. Difficulty thresholds are defined from the distribution of average cloud cover across the three input frames. Specifically, the 90th percentile (top 10%) corresponds to 0.99 average cloud cover, the 80th percentile (top 20%) to 0.78, and the 50th percentile (top 50%) to 0.49. For each sample, cloud cover is computed as the mean fraction of cloudy pixels across the three timestamps. Cloud masks are generated using the s2cloudless algorithm in Google Earth Engine and provided with the dataset.
@@
We use TEC–TT’s prior mean prediction ($\mu$) as a clear-sky estimate of the next frame, capturing large-scale structure while down-weighting transient noise such as clouds (App. F.2.2). Building on this, we adapt TEC–TT for cloud removal by predicting the third input frame from the two previous ones and applying cloud-aware decoding: clear regions retain their original tokens, while cloudy regions are replaced with the transformer’s predicted tokens. Cloud and cloud-shadow masks are available and are downscaled to latent resolution using average pooling.
@@
To provide the clean temporal context, we apply a context reordering heuristic that duplicates the least-cloudy inputs when few clean frames are available. We evaluate 16 variants spanning mask type, mask threshold (0.0 vs. 0.5), context reordering (on vs. off), and decoding policy: Interleave, which predicts only cloudy tokens, and Propagate, which predicts from the first cloudy token onward, plus a pure forecasting baseline that replaces all tokens with predictions. The best setting uses cloud+shadow masks at threshold 0.0, average-pool downscaling, context reordering, and interleave decoding; all cloud-aware variants outperform pure forecasting, which tends to follow seasonal drift rather than reconstruct the target. Intuitively, conditioning only cloudy regions on predictions while preserving clear tokens reduces seasonal bias and allows TEC–TT to focus its temporal prior on reconstructing the true clear-sky target. Throughout the paper (Sec. 5.3, App. G.1), we report results using this best configuration.
@@
Although our introduced codecs are trained on L2A reflectance data, the AllClear benchmark is defined on L1C inputs. To ensure a fair zero-shot evaluation, we therefore pretrained a TEC–TT model on the SSL4EO-S12 v1.1 L1C data modality using a high $\lambda=700$ to focus on reconstruction quality. This model uses the same architecture, hyperparameters, and training procedure as the corresponding L2A variant, and is applied without any task-specific fine-tuning on AllClear.
G.1 Quantitative results

### v2

@@ Appendix 0.C Latent Repacking and FlexTEC
Scope and notation. FlexTEC is the flexible–rate variant of TECTT, obtained by integrating Latent Repacking and masked training. It uses the same analysis/synthesis (ELIC) backbone and temporal transformer stack as TECTT, with latent channel width $d_{\text{lat}}{=}192$. After tokenizing the current frame into T=16T{=}16 tokens per patch, repacking groups channels into $T$ channel-slice tokens. Consequently, FlexTEC exposes 16 discrete quality levels via the token budget $K\in\{1,\dots,16\}$, applied consistently across all patches in an image and frames in a sequence. Each token carries $k=d_{\text{lat}}/T=12$ channels shared across all spatial positions. For all rate–distortion (RD) visualizations in this paper, we report curves for budgets $K=\{1,2,3,4,5,6,7,8,12,16\}$, spanning from the lowest-rate (K=1K{=}1) to the highest-quality (K=16K{=}16) operating points.
Implementation differences vs. TEC-TT. FlexTEC is architecturally identical to TECTT except for: (i) the permutation that repacks tokens (Sec. 3.2); (ii) token masking with a learned mask token $m\in\mathbb{R}^{d_{\text{lat}}}$ used during training to replace dropped tokens; and (iii) the masked-rate objective with budget sampling. All other layers, dimensions, and hyperparameters remain unchanged.
FlexTEC is trained with the same hyperparameters as TECTT, but for 400k steps (vs. 300k for TECTT) to account for the task’s added complexity. A single checkpoint trained at $\lambda{=}800$ (slightly higher than the TECTT maximum of 700) is used to cover the full bitrate range under masked training. Empirically, the $T/K$ scaling in the rate term shifts the effective operating point toward lower rates for the same $\lambda$, motivating this increase.
Objective and inference. With mask $M$, the rate is computed on unmasked tokens and, to prevent information collapse into the earliest tokens, upweighted by $T/K$:
@@
Masking for variable-rate robustness. Repacking latents alone is insufficient: a model trained only with full-token inputs learns to rely on all tokens and collapses when some are dropped. We therefore train with masked budgets: sample $K\in\{1,\dots,T\}$, replace the last TKT{-}K tokens by a learned mask vector $m\in\mathbb{R}^{d_{lat}}$, and compute the rate only on unmasked tokens $R(M)$ (as defined in Eq. 3). For stability we use teacher forcing: masking applies only to the current frame’s tokens for the rate term, while the temporal encoder always consumes the real quantized past latents $(\hat{\mathbf{y}}_{i-2},\hat{\mathbf{y}}_{i-1})$. Inspired by FlexTok [4], budgets are drawn from a categorical distribution biased toward larger values, $\Pr(K{=}k)\propto k$ (i.e., the multiset $\{1,2,2,\dots,T,\dots,T\}$), which trains the model frequently near high-rate operation while still exposing it to low-budget regimes. Masked tokens are replaced by a learnable per-channel vector $m\in\mathbb{R}^{d_{lat}}$ ($d_{lat}{=}192$), initialized uniformly in [1,1][-1,1], and shared across positions. This provides a stable placeholder during training while allowing the image decoder to learn how to interpret missing content.
@@
Inference filling. At inference, dropped tokens are not transmitted. By default, we fill them with the transformer’s predicted means, i.e., $\hat{\mathbf{y}}_{i,b,t}\leftarrow\mu_{i,b,t}$ for frame $i$, latent block $b$, and token $t$, which leverages the learned temporal prior to improve reconstruction quality at a given bitrate. As a lighter alternative (reduced compute), we can instead substitute the learned mask vector $m$ for all dropped tokens.
Appendix 0.D Evaluation details and baseline methods
Baseline quality settings. We evaluate classical codecs across the following quality grids:
@@ Appendix 0.D Evaluation details and baseline methods
Implementations and bit depth. JPEG and WebP are executed via Pillow, supporting only 8-bit input. JPEG2000 is run per band using Glymur/OpenJPEG, which allows for 8- or 16-bit quantization and lossy compression via target ratios. HEVC encoding is performed with FFmpeg and x265, using raw 12-bit monochrome input (pixel format yuv400p12le), CRF=Q=Q, -preset medium, and -tune psnr. Bitstream sizes are measured directly from encoded outputs to compute rate. All codecs operate per band (grayscale), and we report bits-per-pixel-band-frame (bppbf) by aggregating bitstream sizes across all bands and frames.
Rate metric (bppbf). We report rate as bits per pixel–band–frame (bppbf):
@@
Appendix 0.E Quantitative results
@@
0.E.1 PSNR range sensitivity
@@
0.E.2 Temporal conditioning effects
We evaluate how temporal context influences fixed-rate TECTT models at inference. To isolate genuine temporal gains from the overhead of early bootstrap frames, we vary the number of available past frames and additionally analyze P-frame performance.
@@ 0.E.2 Temporal conditioning effects
Figure 12 shows corresponding RD curves evaluated on P-frames only—the last two frames in each four-frame sequence, consistent with TECTT’s training context of two past frames. For non-temporal codecs, this selection simply aligns the evaluation set with TECTT. The performance gap between TECTT and image-only codecs (TECFP, TECELIC) widens under this evaluation, reflecting the amortized cost of the initial bootstrap frames. When temporal conditioning is disabled and TECTT is run in “image mode,” its performance closely follows TECELIC, consistent with their shared analysis/synthesis backbone.
Table 6: Effect of inference context $c$ on TEC-TT (trained with 2-frame context, $\lambda{=}5$). We report bits-per-pixel–band–frame (bppbf, $\downarrow$) and PSNR-per-band (65k, $\uparrow$). Context $c\in\{0,1,1{+}1,2\}$ denotes the number of conditioned frames. The P-frames row evaluates only the last two frames (full context) of each 4-frame sequence. Percent changes are relative to c=0c{=}0.
@@
0.E.3 Flexible-rate behavior
@@
Fig. 13 shows the effect of masking in flexible‐rate training by comparing FlexTEC (Latent Repacking with masking) to a variant trained without masking. The latter deteriorates sharply when tokens are dropped at test time—its RD curve is unstable and substantially below FlexTEC—whereas FlexTEC degrades smoothly and remains roughly parallel to fixed‐rate baselines. This confirms that masking is essential for stable variable‐rate performance.
Fig. 14 compares FlexTEC on P-frames (last two frames, full context under c=2c{=}2) versus all frames. The gap is notably larger than for fixed-rate TEC-TT, particularly at low rates. We hypothesize two compounding effects: (i) latent repacking with masked training encourages FlexTEC to concentrate scene-wide, high-utility content into the earliest tokens of bootstrap frames, lowering their cost relative to fixed-rate models; and (ii) for fully conditioned P-frames, the same curriculum distributes information more evenly across tokens, so truncation retains most of the essentials. Together, these effects yield stronger amortization when excluding bootstrap frames, with the benefit most pronounced in the high-compression regime.
@@
Appendix 0.F Qualitative examples
This section complements the main paper with additional visual examples. We first compare reconstructions across all codecs on representative SSL4EO-S12 samples (Sec. 0.F.1). We then examine model beliefs and forecasting behavior (Sec. 0.F.2) for TEC-TT and FlexTEC. We assesss how dropping tokens impacts TEC-TT and the importance of token masking for Latent Repacking. We also show how TEC–TT forecasts the next frame from past context, contrasting mean predictions with stochastic samples.
@@ Appendix 0.F Qualitative examples
0.F.1 General reconstructions
We compare TEC-TT, TEC-FP, and classical codecs (JPEG2000, WebP) on SSL4EO-S12 v1.1 validation samples at matched rate $\approx 0.20$ bppbf. Each row in Fig. 0.F.1 shows the original and reconstructions from each codec, annotated with the average bppbf and PSNR-per-band (65k clipped) across the sequence.
0.F.2 Model beliefs and forecasting
@@
0.F.2.1 Token budget comparison
Figure 16: Effect of filling dropped tokens (K=5K{=}5) with model prior predictions at inference for different TEC-$T$ variants. Vanilla TEC-TT exhibits spatial holes and banding when tokens are removed. Adding Latent Repacking without masking improves quality but leaves uneven detail, while FlexTEC preserves global layout and reduces artifacts.
@@ 0.F.2.1 Token budget comparison
0.F.2.2 Forecasts from past context
@@
Figure 17: TEC-$T$ forecasts using only past context ($K=0$). Model beliefs are visualized via the mean prediction ($\mu$) and full prior sampling from $\mathcal{N}(\mu,\sigma^{2})$. The $\mu$-forecast reconstructs coherent large-scale structure, while sampling reveals plausible variations (e.g., clouds, surface texture) that reflect the model’s uncertainty.
@@
We probe TEC-TT’s model beliefs by predicting the current frame from past context only, using either the predicted mean or samples from the distribution (Fig. 17). The $\mu$-forecast reliably captures large-scale structure, while sampling from the full prior (mean and variance) expresses context-aware uncertainty, primarily reflecting cloud variability. This illustrates the learned distribution rather than a single point estimate. The conservative single-point forecast ($\mu$-forecast) produces a clear-sky prediction for the next frame, enabling TEC-TT to perform cloud removal as evaluated in AllClear.
Appendix 0.G Additional details on cloud inpainting
@@
We provide extended results for the AllClear cloud inpainting benchmark [66], complementing Sec. 5.3. TEC-TT is applied without task-specific fine-tuning, leveraging only its latent temporal predictions.
@@
Experimental setup. We follow the AllClear evaluation protocol, reporting metrics across all spectral bands (not per-band). We compare against heuristic baselines (LeastCloudy, Mosaicing) and zero-shot neural methods (CTGAN, DiffCR, PMAA, U-TILISE, UnCRtainTS). LeastCloudy selects the input image with the lowest cloud+shadow coverage, while Mosaicing fills each pixel by copying a single clear value, averaging if multiple are clear, or using 0.5 if none are clear. On AllClear, these heuristics rank among the top three zero-shot methods, outperforming most neural baselines without fine-tuning.
@@
Besides reporting metrics on the full test set, we also evaluate subsets stratified by cloudiness. Difficulty thresholds are defined from the distribution of average cloud cover across the three input frames. Specifically, the 90th percentile (top 10%) corresponds to 0.99 average cloud cover, the 80th percentile (top 20%) to 0.78, and the 50th percentile (top 50%) to 0.49. For each sample, cloud cover is computed as the mean fraction of cloudy pixels across the three timestamps. Cloud masks are generated using the s2cloudless algorithm in Google Earth Engine and are provided with the dataset.
@@
We use TEC-TT’s prior mean prediction ($\mu$) as a clear-sky estimate of the next frame, capturing large-scale structure while down-weighting transient noise such as clouds (Sec. 0.F.2.2). Building on this, we adapt TEC-TT for cloud removal by predicting the third input frame from the two previous ones and applying cloud-aware decoding: clear regions retain their original tokens, while cloudy regions are replaced with the transformer’s predicted tokens. Cloud and cloud-shadow masks are available and are downscaled to latent resolution using average pooling.
@@
To provide the clean temporal context, we apply a context reordering heuristic that duplicates the least-cloudy inputs when few clean frames are available. We evaluate 16 variants spanning mask type, mask threshold (0.0 vs. 0.5), context reordering (on vs. off), and decoding policy: Interleave, which predicts only cloudy tokens, and Propagate, which predicts from the first cloudy token onward, plus a pure forecasting baseline that replaces all tokens with predictions. The best setting uses cloud+shadow masks at threshold 0.0, average-pool downscaling, context reordering, and interleave decoding; all cloud-aware variants outperform pure forecasting, which tends to follow seasonal drift rather than reconstruct the target. Intuitively, conditioning only cloudy regions on predictions while preserving clear tokens reduces seasonal bias and allows TEC-TT to focus its temporal prior on reconstructing the true clear-sky target. Throughout the paper (Sec. 5.3 , Sec. 0.G.1), we report results using this best configuration.
@@
Although our introduced codecs are trained on L2A reflectance data, the AllClear benchmark is defined on L1C inputs. To ensure a fair zero-shot evaluation, we therefore pretrained a TEC-TT model on the SSL4EO-S12 v1.1 L1C data modality using a high $\lambda=700$ to focus on reconstruction quality. This model uses the same architecture, hyperparameters, and training procedure as the corresponding L2A variant, and is applied without any task-specific fine-tuning on AllClear.
0.G.1 Quantitative results